# Eye of the vortex: bound spectra in tunable horizonless rotational analogs

### v2

@@ IV.2 Bound states
Figure 1 presents the theoretical predictions of the minimum frequency permissible for propagation of BSs for a given circulation $C$, and a variety of azimuthal numbers $m$ and overtones $n$. We consider both co-rotating ($m>0$) and counter-rotating ($m<0$) BS solutions, from which the symmetry $\omega_{mn}\rightarrow-\omega_{-mn}$ arises. We note that this symmetry holds in our analysis due to the idealization of the experiment, that lacks nonlinear dispersion relations. This symmetry holds for all overtone number $n$. Notice that in some cases of boundary conditions, the discrete BS frequency changes from positive to negative, though since they are purely real, the periodicity of the solution is still preserved. By comparing our resulting spectra with the experimental observations in [68] (see Figs. 2f and 2g), we find that their behavior for different boundary conditions remains qualitatively the same with those found in [68], especially for radii of the superfluid interface that are much larger than the experimental potential barrier maximum. In fact the larger the radius, the more the aforementioned symmetry approximately holds in this particular experiment due to the reduction of the angular velocity distribution. Of course, in our case the choice of $C$ points to a velocity distribution, therefore when choosing $C$, we effectively change the radius of the interface, while ignoring other effects, such as dissipation of the wave’s energy by viscous damping and by scattering into the vortex. Nevertheless, this is somewhat balanced by the stochastic drive originating from the fluid flow and mechanical vibrations in the experimental setup [68], thus our assumptions and resulting spectral space $(m,f)$ of discrete BSs are not necessarily purely theoretical. Another similarity that appears between our data in Fig. 1 and those in Fig. 2 of [68] is the fact that only a certain region of the spectral space is populated with BSs. We, also, observe that only some high-frequency (equivalent to high-energy) waves have the capability to propagate.
@@
Figure 2 shows the BS eigenfunctions with respect to the radial coordinate $r$. The demonstration unveils the pattern of the BSs with respect to the radius of the idealized vortex fluid flow, for a fixed circulation $C=1$. The shapes of the effective potentials, for different overtone numbers $n$, are more pronounced closer to their respective peaks. Different boundary conditions give rise to different BS eigenfunctions, and respective BS frequencies, as well as higher- or lower-frequency maxima of the effective potential peaks. The most pronounced peak appears for the boundary condition (V).
@@
Interestingly, similar co-rotating BSs (dubbed QBSs) have been found in the laboratory experiment involving a giant quantum vortex of superfluid 4He [68]. Even though there is a resemblance between our results and those extracted from the realistic experiment, this is only a qualitative resemblance. Our effective spacetime metric does capture the important qualities and features of the realistic experiment, but it is not yet suitable to draw solid quantitative conclusions, especially because the patterns that appear in actual superfluid vortex experiments, as in [68], regard interface waves, and not acoustic waves as the idealized manifestation of the effective metric and analysis involves. Our results, therefore, are qualitatively robust and may be of significance in the future when all aspects of the experimental setup are taken into account, such as nonlinear dispersion, interface waves, the transition between solid and hollow core vortices and, finally, which choice of boundary conditions is the most suitable one for such an experiment. Our analysis has also been performed with respect to the Mathieu-space radial coordinate $x$, in order to better understand if the boundary conditions of the BS eigenfunctions are met.
@@
Figure 3 demonstrates the absolute values of BS eigenfunctions for the six pairs of boundary conditions with respect to the coordinate $x$. We focus on the co-rotating $m=1$ mode, with circulation set to $C=1$, though these choices are practically random since qualitatively similar results appear for higher $m$ and varying $C$. We observe that in the $x$-space, all the boundary conditions, that are imposed in order to obtain the BS solutions, are appropriately satisfied.
@@
The most important question then is which boundary condition is more physical for the calculation of BSs in their domain of existence. To find QBSs of BHs, the boundary conditions imposed are purely ingoing waves at the BH event horizon and purely reflecting waves at infinity [76]. Here, the experiment we are trying to reproduce possesses no radial inflow of perturbations into the vortex, therefore there is no acoustic event horizon such as those appearing in BHs and BH analogs. Thus, a different boundary condition has to be imposed at the “center” of the vortex $x=0$. Since there is no event horizon, there is also an absence of a singularity at the internal vicinity of the vortex. This leads to the typical boundary condition that is imposed close to the center of stars when their perturbations are studied, i.e. regularity at the central region of the star [77, 78, 79, 80, 81, 82]. The boundary conditions at the center ensure that solutions remain finite and physically plausible. This is also the mathematical requirement to avoid singularities.
@@
One of the most general case of boundary conditions that ensures regularity is the Neumann boundary condition that specifies the values of the derivative applied at the boundary of a domain, which in our case translates to $R^{\prime}(x=0)=\textrm{const.}$. in order for the solutions to remain finite there. This means that there is no flow across the boundary, effectively making the central region a no-flux zone for the perturbation. In this context, it essentially implies that the perturbation is symmetric, as it occurs with solutions to the Mathieu equation which have certain periodicity. The normal derivative of the perturbed quantity (say the temperature or the density of the superfluid in [68]) is set to zero or a constant, meaning there is no change in the slope of the quantity across the internal region.
@@
There are more restrictive boundary conditions, such as the Dirichlet one where the field satisfies $R(x=0)=\textrm{const.}$. at the central region, that is there is no perturbation at the vicinity of the core. Nevertheless, the constant resulting from the Neumann boundary condition can be chosen to be zero or otherwise, in order to cover wider scenarios of regularity, such as regularity at the inner (close to the central) region’s boundary and the outer boundary of the rotational fluid that is deemed to possess only an azimuthal velocity distribution. From a mathematical perspective, the boundary conditions that we can apply in Mathieu differential equations either set the field or its normal derivative equal to zero at $x=0$ [see Eqs. (35)-(40)]. In principle, both a Dirichlet or a Neumann boundary condition can be used here, since they are both physically viable. Nevertheless, we could argue that in order to keep the temperature or the density of the superfluid constant at the central region of the vortex (and not zero, according to [68]), then the most general boundary condition is the Neumann one that ensures both regularity and a non-zero finite constant around $x=0$. From our perspective, this narrows down the possible boundary conditions of the Mathieu equation to (36), (39) and (40).
@@
As mentioned above, to find QBSs (or BSs) in BHs (or horizonless compact objects) we impose purely reflecting boundary conditions at infinity. This corresponds to Dirchlet boundary conditions at infinity, and in our case the eigenfunction should become zero at the boundary of the Mathieu space that our solution exists. Following our argumentation, the pair of boundary conditions needed in order for the resulting BSs to have a more appropriate physical meaning are the ones described by (39). This coincides with the boundary condition (V) that allows for the highest frequency peak of the effective potential, which coincidentally looks more similar to what was found in the experimental apparatus of Ref. [68]. This, of course, is not a statement of quantitative agreement between the metric’s perturbative behavior and the experimental data, but rather a mere coincidence. In principle, all boundary conditions of the Mathieu equation can qualitatively reproduce the shapes and overall structure shown in Figs. 2, 4, and 5 of [68].
@@ V Final remarks
We have proposed the use of an acoustic metric that describes a draining and rotational fluid flow, specified by the drain $D$ and the circulation $C$ parameters included in the metric, in order to qualitatively reproduce current experimental data. This acoustic spacetime has been constructed assuming the hydrodynamic limit, that discards any nonlinear dispersion relations between the velocity density profile and the perturbations that we study. There are an abundance of applications of this metric in order to mimic analog BH experiments. A special case of interest in this work is the giant quantum vortex, recently constructed in the laboratory [68]. This vortex has two configurations; solid or hollow core. Even though the utilized metric does not directly discusses the possibility of two different cores, the BS analysis of acoustic waves shows a lot of qualitative similarities with the actual experiment and seems universal for BSs, irregardless of the core structure, since there is no radial flow of perturbations in the experiment and in the effective metric ($D=0$). By obtaining a wave equation for the propagation of acoustic waves in the analog spacetime, we find a transformation that turns it into a Mathieu differential equation. Therefore, we find the BS solutions and frequencies through the Mathieu functions by applying all possible boundary conditions that these functions must, mathematically, satisfy. We then discuss their structure and compare them with the quasi-periodic interface BSs found in the vortex experiment. Our findings indicate that the phenomenology of such an intricate experiment could be qualitatively reproduced by our extremely simplified metric and perturbation analysis.
We believe that our results for the BSs of acoustic waves might be helpful to the experimental community, since they can describe periodic acoustic waves propagating in a rotational fluid flow without radial inflow. In this sense, we expect that one could test the spacetime background under consideration and hence determine, in principle, if and how the idealized acoustic metric used here can provide more than just qualitative data.
@@
To that end, we need to consider a key ingredient that is not directly discussed in the horizonless tunable acoustic rotational analog. This is the ability to also tune the horizonless metric between a solid and a hollow core vortex. In order to be able to perform such a change to the metric, we need to include a new parameter and/or a functional of an existing parameter that enables us to perform this transition. The effective analog metric includes, besides the drain parameter $D$, a parameter for the circulation $C$ of the vortex. $C$ in turn, traces back to an angular velocity profile and a corresponding propeller frequency that controls the absence or presence of a hollow core. Thus, it might be the case that the circulation can be included in the metric in a form of a functional that enables this transition, depending on its value. This work needs to take advantage of the experimental data from Ref. [68], in order to build (fit) a relation between various $C$ and the corresponding propeller frequencies of the experimental apparatus. We leave this work for the future.
@@
Another interesting feature that can emerge from such experiments, and the resulting idealized metric under consideration, is the fact that there is no inflow of radial perturbations into the core. In a sense, the experiment may serve as a proxy for astrophysical environments; either extremely massive ones that surround galaxies, where the inflow of matter is negligible with respect to the lifetime of the environment (such as dark matter halos) or with dispersion relations that describe high-energy trans-Planckian effects around BHs. Hence, our analysis, and the successful control of the experiment in [68], may lead the way for a more simplistic, yet accurate enough, design of decoupling environmental effects (matter perturbations) from gravitational ones (especially polar gravitational perturbations in astrophysical BHs), and in general the relativistic treatment of astrophysical environments around BHs [83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 90, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109].
@@ Acknowledgements.
The authors are indebted to Prof. Kostas D. Kokkotas for fruitful discussions and comments in the manuscript. The authors also would like to thank Pietro Smaniotto and Leonardo Solidoro for their help in better understanding the experimental apparatus and its function. This study was financed in part by the Conselho Nacional de Desenvolvimento Científico e Tecnológico – Brasil (CNPq) – Research Project No. 440846/2023-4 and Research Fellowship No. 201221/2024-1. H.S.V. is partially supported by the Alexander von Humboldt-Stiftung/Foundation (Grant No. 1209836). Funded by the Federal Ministry of Education and Research (BMBF) and the Baden-Württemberg Ministry of Science as part of the Excellence Strategy of the German Federal and State Governments – Reference No. 1.-31.3.2/0086017037. K.D. acknowledges financial support provided by FCT – Fundação para a Ciência e a Tecnologia, I.P., under the Scientific Employment Stimulus – Individual Call – Grant No. 2023.07417.CEECIND/CP2830/CT0008. This project has received funding from the European Union’s Horizon-MSCA-2022 research and innovation programme “Einstein Waves” under grant agreement No. 101131233.

### v3

@@ IV.2 Bound spectra
Figure 1 presents the theoretical predictions of the minimum frequency permissible for propagation of BSs for a given circulation $C$, and a variety of azimuthal numbers $m$ and overtones $n$. We consider both co-rotating ($m>0$) and counter-rotating ($m<0$) BS eigensolutions, from which the symmetry $\omega_{mn}\rightarrow-\omega_{-mn}$ arises. We note that this symmetry holds in our analysis due to the idealization of the experiment, that lacks nonlinear dispersion relations. This symmetry holds for all overtone number $n$. Notice that in some cases of boundary value conditions, the discrete BS frequency changes from positive to negative, though since they are purely real, the symmetry of the solution is still preserved. By comparing our resulting spectra with the experimental observations in [70] (see Figs. 2f and 2g), we find that their behavior for different boundary conditions remains qualitatively the same with those found in [70], especially for radii of the superfluid interface that are much larger than the experimental potential barrier maximum. In fact the larger the radius, the more the aforementioned symmetry approximately holds in this particular experiment due to the reduction of the angular velocity distribution. Of course, in our case the choice of $C$ points to a velocity distribution, therefore when choosing $C$, we effectively change the radius of the interface, while ignoring other effects, such as dissipation of the wave’s energy by viscous damping and by scattering into the vortex. Nevertheless, this is somewhat balanced by the stochastic drive originating from the fluid flow and mechanical vibrations in the experimental setup [70], thus our assumptions and resulting spectral space $(m,f)$ of discrete BSs are not necessarily purely theoretical. Another similarity that appears between our data in Fig. 1 and those in Fig. 2 of [70] is the fact that only a certain region of the spectral space is populated with BSs. We, also, observe that only some high-frequency (equivalent to high-energy) BSs have the capability to propagate.
@@
Figure 2 shows the BS eigenfunctions with respect to the radial coordinate $r$. The demonstration unveils the pattern of the BSs with respect to the radius of the idealized vortex rotational fluid flow, for a fixed circulation $C=1$. The shapes of the effective potentials, for different overtone numbers $n$, are more pronounced closer to their respective peaks. Different boundary value problems give rise to different BS eigenfunctions, and respective BS eigenfrequencies, as well as higher- or lower-frequency maxima of the effective potential peaks. The most pronounced peak appears for the boundary boundary value problem (V).
@@
Interestingly, similar co-rotating BSs (dubbed QBSs) have been found in the laboratory experiment involving a giant quantum vortex of superfluid 4He [70]. Even though there is a resemblance between our results and those extracted from the realistic experiment, this is purely qualitative. Our effective spacetime metric does capture the important qualities and features of the realistic experiment, but it is not yet suitable to draw solid quantitative conclusions, especially because the patterns that appear in actual superfluid vortex experiments, as in [70], regard interface waves, and not acoustic waves as the idealized manifestation of the effective metric and analysis involved. Our results, therefore, are qualitatively robust and may be of significance in the future when all aspects of the experimental setup are taken into account, such as nonlinear dispersion, interface waves, the transition between solid and hollow core vortices and, finally, which choice of boundary value problem is the most suitable one for such an experiment.
@@
For completeness, we have also analyzed the same problem with respect to the modified Mathieu-space radial coordinate $x$. Figure 3 demonstrates the absolute values of BS eigenfunctions for the six pairs of boundary value conditions with respect to the coordinate $x$. We focus on the co-rotating $m=1$ mode, with circulation set to $C=1$, though these choices are practically random since qualitatively similar results appear for higher $m$ and varying $C$. We observe that in the $x$-space, the boundary value conditions associated with the modified Mathieu equation, in conjunction with the standard Mathieu equation, that are imposed in order to obtain the BS eigensolutions, ensure boundedness and regularity. The cases in Fig. 3, are divided into eigensolutions that start at $R(0)=0$, and those that start at $R(0)=1$. Their asymptotic behavior shows oscillatory decay in all cases. Thus, all wavefunctions are both regular at the core and bounded asymptotically.
@@
The most important question then is which boundary value condition is more physical for the calculation of BSs in their domain of existence. To find QBSs of BHs, the boundary conditions imposed are purely ingoing waves at the BH event horizon and purely reflecting waves at infinity [83]. Here, the experiment we are trying to reproduce possesses no radial inflow of perturbations into the vortex, therefore there is no acoustic event horizon such as those appearing in BHs and BH analogs. Thus, a different boundary condition has to be imposed at the “center” of the vortex $x=0$. Since there is no event horizon, there is also an absence of a singularity at the internal vicinity of the vortex. This leads to the typical boundary condition that is imposed close to the center of stars when their perturbations are studied, i.e. regularity at the center of the star [84, 85, 86, 87, 88, 89]. The boundary condition at the center ensure that solutions remain finite and physically plausible. This is also the mathematical requirement to avoid singularities.
@@
To ensure regularity, the Neumann boundary condition specifies the values of the derivative applied at the boundary of a domain, which in our case may translates to $R^{\prime}(x=0)=\textrm{const.}$ in order for the eigensolutions to remain finite there. This means that there is no flow across the boundary, effectively making the central region a no-flux zone for the perturbation. In this context, it essentially implies that the perturbation is symmetric, as it occurs with solutions to the modified Mathieu equation which have certain characteristic eigenvalues. The normal derivative of the perturbed quantity (say the temperature or the density of the superfluid in [70]) is set to zero or a constant, meaning there is no change in the slope of the quantity across the internal region.
@@
The Dirichlet boundary on the other hand, where $R(x=0)=\textrm{const.}$ at the central region, ensures that there is no perturbation at the vicinity of the core whatsoever. Nevertheless, the constant resulting from the Neumann boundary condition can be chosen to be zero or otherwise, in order to cover wider scenarios of regularity, such as regularity at the inner (close to the central) region’s boundary and the outer boundary of the rotational fluid that is deemed to possess only an azimuthal velocity distribution. From a mathematical perspective, the boundary value conditions that we can apply in modified Mathieu differential equations either set the field or its normal derivative equal to zero at $x=0$ [see Eqs. (43)-(48)]. In principle, both a Dirichlet or a Neumann boundary condition can be used here, since they are both physically viable. Nevertheless, we could argue that in order to keep the temperature or the density of the superfluid constant at the central region of the vortex (and not zero, in accordance with Ref. [70]), then the most general boundary condition is the Neumann one that ensures both regularity and a non-zero finite constant at $x=0$. From our perspective, this narrows down the possible boundary value conditions of the modified Mathieu equation to (44), (47) and (48).
@@
As mentioned above, to find QBSs (or BSs) in BHs (or horizonless compact objects) we impose purely reflecting boundary conditions at infinity. This corresponds to Dirichlet boundary conditions at infinity, and in our case the eigenfunction should become zero at the asymptotic boundary of the modified Mathieu space that our solution exists. Following our argumentation, the pair of boundary value conditions needed in order for the resulting BSs to have a more appropriate physical meaning are the ones described by Eq. (39). This coincides with the boundary value condition (V) that allows for the highest frequency peak of the effective potential, which coincidentally looks more similar to what was found in the experimental apparatus of Ref. [70]. This, of course, is not a statement of quantitative agreement between the metric’s perturbative behavior and the experimental data, but rather a mere coincidence. In principle, all boundary value conditions of the modified Mathieu equation can qualitatively reproduce the shapes and overall structure shown in Figs. 2, 4, and 5 of [70].
@@ V Final remarks
We have proposed the use of an acoustic metric that describes a draining and rotational fluid flow, specified by the drain $D$ and the circulation $C$ parameters, in order to qualitatively reproduce recent patterns of experimental data. This acoustic spacetime has been constructed assuming the hydrodynamic limit, that discards any nonlinear dispersion relations between the velocity density profile and the perturbations that we study. There are an abundance of applications of this metric in order to mimic analog BH experiments. A special case of interest in this work is the giant quantum vortex, recently constructed in the laboratory [70, 71, 72]. The BS analysis of acoustic waves shows a lot of qualitative similarities with the actual experiment and seems universal, since there is no radial flow of perturbations in the experiment and in the effective metric ($D=0$). By obtaining a wave equation for the propagation of acoustic waves in the analog spacetime, we find a transformation that turns it into a modified Mathieu differential equation. Therefore, we find the solutions and frequencies through the modified Mathieu functions by solving all possible boundary value problems that these functions must, mathematically, satisfy. We then discuss their structure and compare them with the quasi-periodic interface BSs found in vortex experiments. Our findings indicate that the phenomenology of such an experiment could be qualitatively reproduced by our extremely simplified metric and perturbation analysis. We believe that our results for the BSs of acoustic waves might be helpful to the experimental community, since they can describe bounded acoustic waves propagating in a rotational fluid flow without radial inflow. In this sense, we expect that one could test the spacetime background under consideration and hence determine, in principle, if and how the idealized acoustic metric used here can provide more than just a qualitative picture.
@@
To that end, we need to consider a key ingredient that is not directly discussed in the horizonless acoustic rotational analog. This is the ability to also tune the metric between a solid and a hollow core vortex, as it is performed in [70], through particular choices of the propeller frequency. In order to be able to perform such a change to the metric, we need to include a new parameter and/or a functional of an existing parameter that enables us to perform this transition. The effective analog metric includes, besides the drain parameter $D$, a parameter for the circulation $C$ of the vortex. $C$ in turn, traces back to an angular velocity profile and a corresponding propeller frequency that controls the absence or presence of a hollow core. Thus, it might be the case that the circulation can be included in the metric in a form of a functional that enables this transition, depending on its value. This work needs to take advantage of the experimental data from Ref. [70], in order to build (fit) a relation between various $C$ and the corresponding propeller frequencies of the experimental apparatus. We leave this work for the future.
@@
Another interesting feature that can emerge from such experiments, and the resulting idealized metric under consideration, is the fact that there is no inflow of radial perturbations into the core. In a sense, the experiment may serve as a proxy for astrophysical environments; either extremely massive ones that surround galaxies, where the inflow of matter is negligible with respect to the lifetime of the environment (such as dark matter halos), or with dispersion relations that describe high-energy trans-Planckian effects around BHs. Hence, our analysis, and the successful control of the experiment in [70], may lead the way for a more simplistic, yet accurate enough, design of decoupling environmental effects (matter perturbations) from the gravitational sector (especially polar gravitational perturbations in astrophysical BHs), and in general the relativistic treatment of astrophysical environments around BHs [90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 97, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119].
@@ Acknowledgements.
The authors are indebted to the anonymous referee for providing extremely insightful comments that led to a much more polished and mathematically sound version of the manuscript. The authors are, also, indebted to Prof. Kostas D. Kokkotas for fruitful discussions and comments in the manuscript. The authors also would like to thank Pietro Smaniotto and Leonardo Solidoro for their help in better understanding the experimental apparatus and its function. This study was financed in part by the Conselho Nacional de Desenvolvimento Científico e Tecnológico – Brasil (CNPq) – Research Project No. 440846/2023-4 and Research Fellowship No. 201221/2024-1. H.S.V. is partially supported by the Alexander von Humboldt-Stiftung/Foundation (Grant No. 1209836). Funded by the Federal Ministry of Education and Research (BMBF) and the Baden-Württemberg Ministry of Science as part of the Excellence Strategy of the German Federal and State Governments – Reference No. 1.-31.3.2/0086017037. K.D. acknowledges financial support provided by FCT – Fundação para a Ciência e a Tecnologia, I.P., under the Scientific Employment Stimulus – Individual Call – Grant No. 2023.07417.CEECIND/CP2830/CT0008. This project has received funding from the European Union’s Horizon-MSCA-2022 research and innovation programme “Einstein Waves” under grant agreement No. 101131233.
Appendix A Standard Mathieu equation and its periodic solutions
The standard form of the Mathieu equation is given by [75]
$R^{\prime\prime}(x)+[a-2q\cos(2x)]R(x)=0,$ (41)
whose general solution is given by
$R(x)$ $=C_{1}\ \mbox{MathieuC}(a,q,x)$
$+C_{2}\ \mbox{MathieuS}(a,q,x),$ (42)
where $C_{1}$ and $C_{2}$ are constants to be determined, MathieuC and MathieuS are the even and odd standard Mathieu functions, respectively. The Mathieu functions become periodic for some characteristic values of $a$ for given $q$. In any half-open interval of length $\pi$ on the real axis, an even Mathieu function of the first kind with $n$ zeros in the interval $0\leq x<\pi$ is denoted by $\mbox{ce}_{n}(x,q)$, and has characteristic values $a_{n}(q)$, with $n=0,1,2,\ldots$, while an odd Mathieu function of the first kind is denoted by $\mbox{se}_{n}(x,q)$, and has characteristic values $b_{n}(q)$, with $n=1,2,\ldots$. From now on, we will write $\mbox{ce}_{n}(x)$, $\mbox{se}_{n}(x)$, $a_{n}$ and $b_{n}$, by omitting $q$. These characteristic values satisfy any of the following pairs of boundary conditions [75]
(I) $\qquad R(0)=R(\pi)=0,\ \mbox{for}\ \mbox{se}_{n}(x)\ \mbox{and}\ b_{n},\ \mbox{with period of}\ \pi;$ (43)
(II) $\qquad R^{\prime}(0)=R^{\prime}(\pi)=0,\ \mbox{for}\ \mbox{ce}_{n}(x)\ \mbox{and}\ a_{n},\ \mbox{with period of}\ \pi.$ (44)
On the other hand, if $R(x)$ is either $\mbox{se}_{n}(x)$ or $\mbox{ce}_{n}(x)$, then $R(x)$ and $R(\pi-x)$ satisfy the same differential equation and the same boundary conditions, and must be constant multiples of each other so that $R(x)$ is either an even or an odd function of $\pi/2-x$. This leads to four additional pairs of boundary conditions [75]:
(III) $\qquad R(0)=R(\pi/2)=0,\ \mbox{for}\ \mbox{se}_{2n+2}(x)\ \mbox{and}\ b_{2n+2},\ \mbox{with period of}\ \pi;$ (45)
(IV) $\qquad R(0)=R^{\prime}(\pi/2)=0,\ \mbox{for}\ \mbox{se}_{2n+1}(x)\ \mbox{and}\ b_{2n+1},\ \mbox{with period of}\ 2\pi;$ (46)
(V) $\qquad R^{\prime}(0)=R(\pi/2)=0,\ \mbox{for}\ \mbox{ce}_{2n+1}(x)\ \mbox{and}\ a_{2n+1},\ \mbox{with period of}\ 2\pi;$ (47)
(VI) $\qquad R^{\prime}(0)=R^{\prime}(\pi/2)=0,\ \mbox{for}\ \mbox{ce}_{2n}(x)\ \mbox{and}\ a_{2n},\ \mbox{with period of}\ \pi.$ (48)
Here, for each $n=0,1,2,\ldots$ there is exactly one characteristic function of each of these four boundary value problems, and $n$ is the number of zeros in the interval $0<x<\pi/2$.